OpenStudy (anonymous):

Find the length of the major axis of the equation x^2 + 4y^2 = 4

OpenStudy (anonymous):

@dusty

OpenStudy (experimentx):

divide by 4 ... what do you get?

OpenStudy (anonymous):

my answer is 2 but idk if its right or wrong

OpenStudy (anonymous):

reduce it to the form of \[x ^{2}/a ^{2} + y ^{2}/b ^{2}= 1\]

OpenStudy (anonymous):

you should get \[x ^{2}/4 + y ^{2}/1 = 1 \]

OpenStudy (anonymous):

length of the major axis is ' 2a ' if a>b and is ' 2b ' if b>a

OpenStudy (anonymous):

so for b i have to put 1

OpenStudy (anonymous):

yes . a^2 = 4 so a=2 and b^2=1 so b=1 .. as a>b , length is 2a

OpenStudy (anonymous):

ohh.. thank you :)

OpenStudy (anonymous):

can you guys help me with another one ?